# FourierKAN-GCF: Fourier Kolmogorov-Arnold Network - An Effective and Efficient Feature Transformation for Graph Collaborative Filtering

### v1

@@ Abstract.
Recommendation, Collaborative Filtering, Graph Neural Network, Kolmogorov-Arnold Network, Fourier Transform
@@ 3.2. Fourier KAN
However, KAN is more difficult to train than MLP due to the spline function, which does not satisfy our motivation to provide a more effective and efficient method to replace $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ part. The essence of KAN is to form an arbitrary function by superposition of multiple nonlinear functions. Our goal can be converted into finding the split from a complex function into multiple relatively simple nonlinear functions. Naturally, the Fourier transform (Nussbaumer and Nussbaumer, 1982) is an interesting choice. Therefore, we propose the following equation:
@@
where $d$ is the dimension number of features. Fourier coefficients $a_{ik}$ and $b_{ik}$ are trainable. Hyper-parameter $g$ is the gridsize, which plays a critical role in controlling the number of terms (frequencies) used in the Fourier series expansion. Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier transform corresponding to each input dimension. The Fourier transform has a significant advantage in computational efficiency and solves the training difficulty caused by the spline function.
@@ 5. Conclusion
In this paper, we revisit the processes of feature transformation and nonlinear operations during message passing in the NGCF. We assert that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN but add complexity in training. To reduce the training difficulty while benefiting from these representational enhancements, we introduce a novel feature transformation method called FourierKAN. Inspired by KAN, FourierKAN employs a Fourier transform instead of the Spline function in standard KAN, which further reduces the difficulty of training while increasing the representation power. Furthermore, based on FourierKAN, we introduce a simple and effective graph-based recommendation model (FourierKAN-GCF) with extra dropout strategies. Extensive experiments on public datasets demonstrate the superiority of our model over the state-of-the-art methods.

### v2

@@ Abstract.
Recommendation, Collaborative Filtering, Graph Neural Network, Kolmogorov-Arnold Network, Fourier Coefficients
@@ 3.2. Fourier KAN
However, KAN is more difficult to train than MLP due to the spline function, which does not satisfy our motivation to provide a more effective and efficient method to replace $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ part. The essence of KAN is to form an arbitrary function by superposition of multiple nonlinear functions. Our goal can be converted into finding the split from a complex function into multiple relatively simple nonlinear functions. Naturally, the Fourier Coefficients (Nussbaumer and Nussbaumer, 1982) is an interesting choice. Therefore, we propose the following equation:
@@
where $d$ is the dimension number of features. Fourier coefficients $a_{ik}$ and $b_{ik}$ are trainable. Hyper-parameter $g$ is the gridsize, which plays a critical role in controlling the number of terms (frequencies) used in the Fourier series expansion. Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier Coefficients corresponding to each input dimension. The Fourier Coefficients has a significant advantage in computational efficiency and solves the training difficulty caused by the spline function.
@@ 5. Conclusion
In this paper, we revisit the processes of feature transformation and nonlinear operations during message passing in the NGCF. We assert that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN but add complexity in training. To reduce the training difficulty while benefiting from these representational enhancements, we introduce a novel feature transformation method called FourierKAN. Inspired by KAN, FourierKAN employs a Fourier Coefficients instead of the Spline function in standard KAN, which further reduces the difficulty of training while increasing the representation power. Furthermore, based on FourierKAN, we introduce a simple and effective graph-based recommendation model (FourierKAN-GCF) with extra dropout strategies. Extensive experiments on public datasets demonstrate the superiority of our model over the state-of-the-art methods.